# A Transfer Attack to Image Watermarks

### v2

@@ Appendix A Proof of Lemma 1
Appendix B Proof of Theorem 1
@@
Appendix C Proof of Lemma 13
@@
Appendix D Proof of Theorem 2
@@
Appendix E Proof of Theorem 15
The conditional expectation of $|T(x_{w}+\delta)_{j}-T(x_{w})_{j}|$ can be represented as:
@@ Appendix E Proof of Theorem 15
According to Assumption 10, the conditional expectation of $|T(x_{w})_{j}-w_{j}|$ can be represented as:
@@
Therefore, based on Equation 19, we have the following:
@@
Table 1: The detailed settings of $\alpha$ for different number of surrogate models.
# of Surrogate Models 1 5 10 20 30 40 60 80 100
$\alpha$ 0.1 1 2 4
(a) 20 bits
(b) 30 bits
(c) 64 bits
Figure 9: Comparing the average perturbations of our transfer attack and common post-processing methods when they achieve the same evasion rate. The target model uses ResNet architecture and different watermark lengths (the three columns). Dataset is Midjourney.
(a) 20 bits
(b) 30 bits
(c) 64 bits
Figure 10: Comparing the average $\ell_{2}$-norm perturbations of our transfer attack and common post-processing methods when they achieve the same evasion rate. The target model uses ResNet architecture and different watermark lengths (the three columns). First row: Stable Diffusion. Second row: Midjourney.
(a) 20 bits
(b) 30 bits
(c) 64 bits
Figure 11: Comparing the average SSIM of our transfer attack and common post-processing methods when they achieve the same evasion rate. The target model uses ResNet architecture and different watermark lengths (the three columns). First row: Stable Diffusion. Second row: Midjourney.
Figure 12: Comparing the evasion rates of existing and our transfer attacks when the target model is ResNet and uses watermarks with different lengths. Dataset is Midjourney.
(a) $\ell_{2}$-norm perturbation
(b) SSIM
Figure 13: Average $\ell_{2}$-norm perturbations and average SSIM of our transfer attack with different $\epsilon$. The target model uses CNN architecture and watermark length of 30 bits.
(a) $\ell_{2}$-norm perturbation
(b) SSIM
Figure 14: Average $\ell_{2}$-norm perturbations and average SSIM of different variants of our transfer attack. The target model uses CNN architecture and watermark length of 30 bits.

### v4

@@ Appendix C Proof of Lemma 1
Appendix D Proof of Theorem 2
@@
Appendix E Proof of Lemma 13
@@
Appendix F Proof of Theorem 3
@@
Appendix G Proof of Theorem 1
When $1\leq j\leq n$, the conditional expectation of $|T(x_{w}+\delta)_{j}-T(x_{w})_{j}|$ can be represented as:
@@ Appendix G Proof of Theorem 1
According to Assumption 10, when $1\leq j\leq n$, the conditional expectation of $|T(x_{w})_{j}-w_{j}|$ can be represented as:
@@
Therefore, based on Equation 18, when $1\leq j\leq n$, we have the following:
@@
If $n_{t}>n$, when $n<j\leq n_{t}$, we have the following:
$0\leq Pr(T(x_{w}+\delta)_{j}=w_{j}\mid S_{1}(x_{w}+\delta)_{j},\cdots,S_{m}(x_{w}+\delta)_{j})\leq 1$
Therefore, we have the following:
[…]
Appendix H Details of datasets
For the target watermarking models, we use two public datasets (Wang et al., 2023; Turc & Nemade, 2022) generated by Stable Diffusion and Midjourney respectively. Following HiDDeN (Zhu et al., 2018), we randomly select 10,000 images from each dataset to train the target watermarking encoders and decoders. For testing, we randomly sample 1,000 images from the testing set of each dataset, embed the ground-truth watermark into each of them using a target encoder, and then find the perturbation to each watermarked image using different methods. To train the surrogate watermarking models, we sample 10,000 images from another public dataset (Images, 2023) generated by DALL-E 2, i.e., the surrogate dataset consists of these 10,000 images. The input image size of the watermarking models is $128\times 128$ pixels.
Appendix I Details of target watermarking models and watermark lengths
For HiDDeN, we consider two architectures for the target watermarking model: CNN and ResNet. For the CNN architecture, the encoder consists of 4 convolutional blocks, while the decoder consists of 7 convolutional blocks. Each block integrates a Convolution layer, Batch Normalization, and ReLU activation. For the ResNet architecture, the encoder consists of 7 convolutional blocks and the decoder is the ResNet-18.
For StegaStamp, we use model architecture introduced by  Tancik et al. (2020). For Stable Signature, we use the public watermarking model provided by  Fernandez et al. (2023) as the target watermarking model. Smoothed HiDDeN and Smoothed StegaStamp are certifiably robust against bounded adversarial perturbations. For them, we adopt the regression smoothing (Jiang et al., 2024) to obtain the smoothed versions of the corresponding HiDDeN and StegaStamp watermarking models.
Appendix J Details of common post-processing methods
Specifically, we consider the following common post-processing methods.
Gaussian blur. This method smooths the image by averaging pixel values with their neighbors. It applies a Gaussian filter with kernel size of $k\times k$ to an image, characterized by a bell-shaped curve with a zero mean and a standard deviation of $\sigma$. A larger $\sigma$ causes more pronounced blurring, which results in lower watermark detection rate and image quality. Following  Jiang et al. (2023), we set $k=5$ and vary $\sigma$.
Brightness/Contrast. This method modifies the brightness and contrast of an image by adjusting pixel values throughout the image. Specifically, it operates by either increasing or decreasing these values to make the image brighter or darker. The process is regulated by two factors: a brightness factor $B$ and a contrast factor $C$. Formally, for each pixel value $v$, the method transforms it to $Cv+B$. Following  Jiang et al. (2023), we set $B=0.2$ and vary $C$.
Appendix K Details of transfer attacks
For the former 5 transfer attacks, WEvade-B-S is watermark-based, while the other four are classifier-based. In particular, MI-CWA leverages the state-of-the-art multiple surrogate classifiers based transfer adversarial examples.
WEvade-B-S (Jiang et al., 2023). This method trains one surrogate watermarking model. A watermarked image is perturbed such that the watermark decoded by the surrogate decoder for the perturbed image matches with a preset random watermark. In our experiments, we train the surrogate watermarking model on 10,000 images generated by DALL-E 2. To give advantages to this transfer attack, we assume the attacker knows the architecture of the target watermarking model and uses it for the surrogate watermarking model.
AdvEmb-RN18 (An et al., 2024). This method uses a ResNet-18 pretrained on ImageNet to generate a feature embedding for a watermarked image. Then it perturbs the image such that its embedding lies far from the one of the watermarked image.
AdvCls-Real&WM (An et al., 2024). This method trains a surrogate classifier using watermarked and non-watermarked images. The watermarked images are generated by the target GenAI and are watermarked by the target watermarking model, and the non-watermarked images are drawn from a distribution different from the one of the images generated by the target GenAI. A watermarked image is perturbed such that it is misclassified by the surrogate classifier as a non-watermarked image. In our experiments, we utilize images generated by DALL-E 2 as the non-watermarked images. More specifically, the training set for the surrogate classifier comprises 8,000 images generated by Stable Diffusion (or Midjourney) and watermarked by the target watermarking model, and 8,000 non-watermarked images generated by DALL-E 2. The surrogate classifier is based on the ResNet-18 architecture.
AdvCls-Enc-WM1&WM2. This method is the same as AdvCls-Real&WM except for the training data used for the surrogate classifier. It assumes that the attacker has access to the target watermarking encoder and can use it to embed any watermark into an image. The surrogate classifier is trained to distinguish the images embedded with the ground-truth watermark and those embedded with an attacker-chosen watermark. In our experiments, the training set for the surrogate classifier comprises 8,000 images generated by the target GenAI and watermarked by the target watermarking model with one watermark, and 8,000 images generated by DALL-E 2 and watermarked by the target watermarking model with another watermark, where both watermarks are randomly picked.
MI-CWA (Chen et al., 2024). Existing classifier-based transfer attacks (i.e., the above three) only leverage one surrogate classifier. We extend state-of-the-art multiple surrogate classifiers based transfer adversarial examples (Chen et al., 2024) to watermarks. Given the same training dataset as AdvCls-Real&WM, we train 100 surrogate classifiers, each of which has the ResNet-18 architecture. Given a watermarked image, this transfer attack perturbs it such that it is misclassified by the surrogate classifiers as non-watermarked.
DiffPure (Nie et al., 2022). This method uses diffusion model to purify adversarial examples. We extend it to remove watermark. Specifically, DiffPure first adds Gaussian noise gradually to turn a watermarked image into a noised image. Then diffusion model is used to predict the noise step by step to get the denoised image. We use its public code.
Appendix L Details of parameter settings
Specifically, for common post-processing methods, we consider the following range of parameters during adversarial training for a target watermarking model: $Q\in[10,99]$ for JPEG, $\sigma\in[0,0.1]$ for Gaussian noise, $\sigma\in[0,2.0]$ for Gaussian blur, and $C\in[1,3]$ for Brightness/Contrast. For surrogate watermarking models, we adopt a smaller range of parameters for some common post-processing methods during adversarial training to achieve weaker robustness than the target watermarking model as follows: $Q\in[50,99]$ for JPEG, $\sigma\in[0,0.1]$ for Gaussian noise, $\sigma\in[0,1.0]$ for Gaussian blur, and $C\in[1,3]$ for Brightness/Contrast.
By default, we set maximum number of iterations $max\_iter=5,000$, perturbation budget $r=0.25$, sensitivity $\epsilon=0.2$, and learning rate $\alpha=0.1$ for our transfer attack. Unless otherwise mentioned, we use Inverse-Decode to select a target watermark for a surrogate decoder, and Ensemble-Optimization to find the perturbation. $\alpha$ is increased when the number of surrogate watermarking models increases in order to satisfy the constraints of our optimization problem within $5,000$ iterations. The detailed settings for $\alpha$ for different number of surrogate models are shown in Table 1 in Appendix. Moreover, we use $\ell_{2}$-distance as the distance metric $l(\cdot,\cdot)$ for two watermarks.